At one time, the population of Petticoat Junction was a perfect square. Later, with an increase of 100, the population was 1 greater than a perfect square. Now, with an additional increase of 100, the population is again a perfect square. What was the original population?
Answer by ikleyn(52781)

.
2401 persons.


49^2 = 2401,

50^2 = 2500,

51^2 = 2601.